Please solve 6.66. Thank you in advanced.

6-66 Letxand y be independent random variables with variances o2 and o2,respectively. Consider the sum z=ax + (1-a)y 0-as l Find a that minimizes the variance of z.
Answer #1

for X and Y are independent

therefore Var(Z) =a2Var(X)+(1-a)2 *Var(Y)

=a2\sigma12+(1-a)2*\sigma22

for above to minimum: (d/da)Var(Z) =0

(d/da)*(a2\sigma12+(1-a)2*\sigma22) =2a\sigma12-2(1-a)*\sigma22 =0

(1-a)/a =\sigma12/\sigma22

1/a-1 =\sigma12/\sigma22

a =\sigma22/(\sigma12+\sigma22)